OpenStudy (anonymous):

Is this Correct?? (6, 3) and (8, 6). Part 1: Slope (m) = change in y over change in x. Part 2: Now find the equation using the point slope formula: Point Slope Formula = y – y1 = m (x – x1) Part 3: Give an example of an equation for a line parallel to your answer from Part 2 Part 4: Give an example of an equation for a line perpendicular to your answer from Part 2

OpenStudy (anonymous):

Part 1: m= 6 - 3/8 - 6 = 3/2 = 1.5 3/2 = m (Slope) Part 2: Point Slope Formula = y – y1 = m (x – x1) y - 3 = 3/2 (x - 6) y - 3 = x + 6 y- 3 +3 = x +6 +3 y = x + 9

OpenStudy (raden):

i starting to repair from y - 3 = 3/2 (x - 6)

OpenStudy (anonymous):

im a lil confused if i got it right cuz for my homework it says for Part 2 Your final answer should be written in slope intercept form. (y = mx + b)

OpenStudy (raden):

y - 3 = 3/2 (x - 6) multiply by 2 to both sides, it will giving us (y-3) * 2 = 3/2 * 2 * (x-6) 2y - 6 = 3(x-6) 2y-6 = 3x-18 add by 6 to both to both sides : 2y-6+6 = 3x-18+6 2y = 3x-12 finally, divide by 2 to both sides : 2y/2 = (3x-12)/2 y = (3/2)x - 6

OpenStudy (anonymous):

where do you get the 2y

OpenStudy (raden):